A line passing through (4, 11) and (1, y) is parallel to a line with slope 2. What is the value of y?
Can someone show me how to work this problem out. Thanks
Answer by nerdybill(7384) 
 A line passing through (4, 11) and (1, y) is parallel to a line with slope 2. What is the value of y?
.
If two lines are parallel, their slopes MUST be the same.
.
We can calculate the slope of a line given two points by using:
slope = (y2-y1)/(x2-x1)
.
if we assign the following:
(x1,y1) is (4, 11)
and
(x2,y2) is (1, y)
.
Now, simply "plug in" the given values and solve for y:
slope = (y2-y1)/(x2-x1)
2 = (y-11)/(1-4)
2 = (y-11)/(-3)
-2/3 = y-11
11-2/3 = y
33/3 - 2/3 = y
31/3 = y
or
10.333 = y
| |
|